If RS=23-2x,ST=9x-5,and RT=39 find RS.
RS=

If LN=6x-5,LM=x+7, and MN=3x+20 find MN
MN

Find X Round tithe nearest hundredth.
X(-9,2) and Y(5,-4)

Find the midpoint of AB if A(-3,8) and B(-7,-6)

If CE=10x+18,DE=7x-1 and BC=9x-2, find AB
X=
AB=
1
Expert's answer
2020-09-24T15:00:17-0400

"RS + ST = RT"

"23-2x+9x-5=39"

"18+7x=39"

"7x=21"

"x=3"

"RS = 23-2*3=23-6=17"



"LN+LM=MN"

"6x-5+7+x=3x+20"

"4x=18"

"x=4.5"

"MN=3*4.5+20=13.5+20=33.5"


Find the midpoint of AB if A(-3,8) and B(-7,-6)

"x_c = \dfrac{x_a+x_b}{2}=\dfrac{-3-7}{2}=-5"

"y_c = \dfrac{y_a+y_b}{2}=\dfrac{8-6}{2}=1"

C(-5;1) - the midpoint of AB


If CE=10x+18,AE=7x-1 and BC=9x-2, find AB

"AB=AE-BC-CE"

"AB = 7x-1-9x+2-10x-18=-12x-17"